Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 326
Entry time: 18 May 2022, 11:02
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 18.05.2022 |
|
|
General Announcement:
- The cold compressor of the accelerator failed yesterday night at around 23:15. The current expectation is that beam shall be back on Thursday morning.
Instruments and DRC were informed by the PRC via phone. A status entry can be found in the TTfinfo e-log, Operation-2022 e-log and XO Alfresco landing page.
- Update on the summer maintenance planning: Some groups have the review process finalized. Scheduling meetings are ongoing.
If a group thinks it is needed to have an scheduling meeting please feel free to contact Martin Doering.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing to report.
- SA3 PP-L parts that were not behaving properly are under investigation with LAS.
VAC:
- Nothing to report.
XPD:
- Nothing to report.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- Not present.
UNSYS:
- High doses observed in Apple-X are under investigation; A plan to move undulators in the last night did not take place.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Yesterday a loop update cost time with loosing motor parameters but solved at the end.
- Broker pileup: root cause was the physical location of the board (configuration) was changed in between times. Location was changed when configuration was requested.
HED:
- Nothing to report.
SCS:
- Fast sample scanner has problems and needs to be fixed. Involved groups are informed.
SQS:
- Nothing to report. |